# GCondenser: Benchmarking Graph Condensation

### v2

@@ 4.1 Baselines
Trajectory matching: SFGC [36] firstly generates training trajectories of training a GNN backbone using the original graph. For the condensation phase, a GNN with the same architecture is trained with the condensed graph, and the model parameters are matched to the generated model parameters.
@@ 4.2 Implementations
Table 2: Overall node classification performance. For each dataset with different condensed graph sizes and backbone GNN models, the red and bold results indicate the best performance, while the blue and italic results indicate the runner-up.
@@
Dataset GNN Size Ratio Traditional Gradient Distribution Trajectory Whole
@@
Random k-Center GCond GCondX DosCond DosCondX SGDD GCDM GCDMX DM SFGC
Transductive CiteSeer SGC 30 0.9% 53.8$\pm$0.1 52.7$\pm$0.0 71.9$\pm$0.6 65.3$\pm$0.1 71.1$\pm$0.9 64.3$\pm$0.4 71.1$\pm$0.1 66.0$\pm$2.2 70.1$\pm$0.8 64.2$\pm$8.5 65.2$\pm$0.3 70.3±1.0
60 1.8% 61.9$\pm$0.0 66.8$\pm$0.0 71.0$\pm$0.6 70.9$\pm$0.3 72.0$\pm$1.1 70.6$\pm$0.1 69.9$\pm$0.1 66.7$\pm$0.0 65.8$\pm$1.3 65.7$\pm$2.5 67.0$\pm$0.8
120 3.6% 68.1$\pm$0.0 68.1$\pm$0.0 72.5$\pm$1.2 71.5$\pm$0.0 70.6$\pm$0.2 71.8$\pm$0.3 70.8$\pm$0.8 69.1$\pm$1.2 68.1$\pm$0.2 69.7$\pm$0.1 68.8$\pm$0.2
GCN 30 0.9% 58.9$\pm$0.0 65.0$\pm$0.0 46.3$\pm$7.0 66.7$\pm$4.2 69.2$\pm$0.5 67.7$\pm$1.2 70.6$\pm$1.5 71.2$\pm$0.8 72.8$\pm$0.3 72.6$\pm$0.6 69.7$\pm$0.3 71.4±0.5
60 1.8% 62.6$\pm$0.0 67.8$\pm$0.0 54.2$\pm$3.9 69.6$\pm$0.4 70.4$\pm$1.7 70.1$\pm$0.2 71.5$\pm$0.7 71.9$\pm$0.7 71.7$\pm$0.2 72.2$\pm$0.7 69.4$\pm$0.0
120 3.6% 69.6$\pm$0.0 69.4$\pm$0.0 70.7$\pm$0.7 70.9$\pm$0.2 47.3$\pm$7.3 70.4$\pm$0.2 71.0$\pm$0.7 72.3$\pm$1.3 72.5$\pm$0.5 72.4$\pm$0.1 69.8$\pm$0.5
Cora SGC 35 1.3% 59.3$\pm$0.0 63.8$\pm$0.0 80.6$\pm$0.1 80.6$\pm$0.2 80.6$\pm$0.1 80.8$\pm$0.1 62.4$\pm$5.5 77.0$\pm$0.4 79.9$\pm$0.1 77.0$\pm$0.5 73.8$\pm$1.5 79.2±0.6
70 2.6% 70.7$\pm$0.0 70.3$\pm$0.0 81.0$\pm$0.2 79.0$\pm$0.3 80.3$\pm$0.5 79.8$\pm$0.2 80.8$\pm$0.4 78.9$\pm$1.0 79.7$\pm$0.3 78.0$\pm$1.6 77.5$\pm$0.1
140 5.2% 77.0$\pm$0.0 77.1$\pm$0.0 80.9$\pm$0.4 81.2$\pm$0.3 80.8$\pm$0.4 80.8$\pm$0.5 81.4$\pm$0.4 77.9$\pm$0.7 80.1$\pm$0.2 79.6$\pm$0.6 79.2$\pm$0.1
GCN 35 1.3% 63.9$\pm$0.1 66.5$\pm$0.0 80.5$\pm$0.4 79.8$\pm$1.4 72.0$\pm$8.7 80.1$\pm$0.9 80.5$\pm$0.4 78.9$\pm$0.8 79.1$\pm$0.9 79.2$\pm$0.2 79.6$\pm$0.2 81.7±0.9
70 2.6% 73.0$\pm$0.0 71.6$\pm$0.0 78.1$\pm$3.6 80.6$\pm$0.9 79.6$\pm$0.7 81.1$\pm$0.3 81.2$\pm$0.6 79.4$\pm$0.6 80.5$\pm$0.3 79.6$\pm$0.3 79.5$\pm$0.1
140 5.2% 77.1$\pm$0.0 76.6$\pm$0.0 80.2$\pm$1.7 81.5$\pm$0.1 80.5$\pm$0.7 80.6$\pm$0.2 79.9$\pm$1.6 79.9$\pm$0.2 80.2$\pm$0.5 79.9$\pm$0.3 80.1$\pm$0.6
PubMed SGC 15 0.08% 67.8$\pm$0.2 70.5$\pm$0.1 75.9$\pm$0.7 77.3$\pm$0.2 74.2$\pm$1.1 75.7$\pm$0.5 76.4$\pm$0.9 73.3$\pm$1.2 74.0$\pm$0.3 72.1$\pm$0.9 73.9$\pm$0.5 76.9±0.1
30 0.15% 72.5$\pm$0.2 75.8$\pm$0.0 75.2$\pm$0.0 77.1$\pm$0.2 77.2$\pm$0.1 77.0$\pm$0.1 78.0$\pm$0.3 74.7$\pm$0.6 75.2$\pm$0.7 75.1$\pm$0.6 75.8$\pm$0.2
60 0.3% 75.6$\pm$0.0 75.7$\pm$0.0 75.7$\pm$0.0 76.8$\pm$0.1 75.5$\pm$0.2 75.5$\pm$0.0 76.1$\pm$0.1 76.5$\pm$1.1 76.3$\pm$0.2 75.8$\pm$0.0 75.8$\pm$0.0
GCN 15 0.08% 69.8$\pm$0.1 72.1$\pm$0.1 67.6$\pm$10.4 77.8$\pm$0.3 75.7$\pm$0.7 75.7$\pm$0.1 76.7$\pm$1.1 75.9$\pm$0.6 77.1$\pm$0.3 76.0$\pm$0.7 78.4$\pm$0.1 79.3±0.3
30 0.15% 73.7$\pm$0.1 76.4$\pm$0.0 74.6$\pm$0.8 78.0$\pm$0.5 76.8$\pm$0.2 78.6$\pm$0.2 78.5$\pm$0.4 77.4$\pm$0.4 76.8$\pm$0.6 77.5$\pm$0.1 78.1$\pm$0.4
60 0.3% 78.0$\pm$0.0 78.2$\pm$0.0 77.2$\pm$0.7 78.0$\pm$0.1 77.3$\pm$1.2 78.0$\pm$0.1 78.0$\pm$1.1 77.6$\pm$0.4 78.1$\pm$0.3 78.0$\pm$0.2 78.5$\pm$0.5
Arxiv SGC 90 0.05% 45.6$\pm$0.2 51.8$\pm$0.2 65.5$\pm$0.0 66.0$\pm$0.2 62.7$\pm$0.6 61.6$\pm$0.3 64.5$\pm$0.9 60.8$\pm$0.1 59.8$\pm$0.3 61.0$\pm$0.2 66.1$\pm$0.2 68.8±0.0
454 0.25% 55.2$\pm$0.0 58.2$\pm$0.0 66.5$\pm$0.5 66.4$\pm$0.1 63.7$\pm$0.2 63.8$\pm$0.1 66.4$\pm$0.3 62.7$\pm$0.9 61.8$\pm$0.5 62.9$\pm$0.2 66.7$\pm$0.3
909 0.5% 58.3$\pm$0.0 60.3$\pm$0.0 67.2$\pm$0.1 67.4$\pm$0.3 63.9$\pm$0.1 64.3$\pm$0.4 66.9$\pm$0.3 62.4$\pm$0.2 62.6$\pm$0.2 62.5$\pm$0.0 66.4$\pm$0.3
GCN 90 0.05% 47.1$\pm$0.0 54.5$\pm$0.0 53.7$\pm$1.6 62.7$\pm$0.4 55.6$\pm$0.4 61.0$\pm$0.5 55.9$\pm$5.8 63.3$\pm$0.3 63.8$\pm$0.3 64.4$\pm$0.5 66.1$\pm$0.4 71.1±0.0
454 0.25% 56.8$\pm$0.0 60.3$\pm$0.0 64.2$\pm$0.2 65.4$\pm$0.4 61.6$\pm$0.5 64.7$\pm$0.2 63.2$\pm$0.3 66.4$\pm$0.1 66.7$\pm$0.4 67.5$\pm$0.3 67.2$\pm$0.4
909 0.5% 60.3$\pm$0.0 62.1$\pm$0.0 65.1$\pm$0.4 66.6$\pm$0.2 63.4$\pm$0.4 65.8$\pm$0.1 66.8$\pm$0.3 67.6$\pm$0.0 67.6$\pm$0.3 68.0$\pm$0.3 67.8$\pm$0.2
Products SGC 612 0.025% 51.6$\pm$1.3 48.6$\pm$0.6 64.0$\pm$0.2 64.6$\pm$0.1 62.1$\pm$0.1 63.6$\pm$0.3 64.9$\pm$0.1 57.7$\pm$0.2 58.9$\pm$0.1 58.5$\pm$0.6 62.2$\pm$0.1 64.7±0.1
1225 0.05% 55.2$\pm$0.8 52.2$\pm$0.7 64.0$\pm$0.1 62.4$\pm$0.1 61.0$\pm$0.3 62.7$\pm$0.4 62.3$\pm$0.2 58.2$\pm$0.3 61.0$\pm$0.1 60.9$\pm$0.1 62.2$\pm$0.2
2449 0.1% 58.0$\pm$0.8 55.4$\pm$0.4 64.4$\pm$0.4 62.8$\pm$0.1 61.4$\pm$0.2 62.3$\pm$0.2 64.3$\pm$0.3 60.8$\pm$0.2 61.3$\pm$0.1 61.3$\pm$0.2 61.9$\pm$0.2
GCN 612 0.025% 57.8$\pm$0.5 55.4$\pm$0.8 63.7$\pm$0.3 65.5$\pm$0.2 61.2$\pm$0.3 62.2$\pm$0.6 64.0$\pm$0.4 66.5$\pm$0.1 68.0$\pm$0.3 66.2$\pm$0.1 67.1$\pm$0.2 73.1±0.1
1225 0.05% 61.6$\pm$0.5 57.6$\pm$0.7 67.0$\pm$0.2 66.6$\pm$0.3 62.6$\pm$0.4 64.2$\pm$0.2 65.9$\pm$0.2 68.4$\pm$0.4 68.8$\pm$0.1 68.5$\pm$0.2 67.9$\pm$0.3
2449 0.1% 65.3$\pm$0.5 59.1$\pm$0.5 68.0$\pm$0.2 68.3$\pm$0.2 65.8$\pm$0.2 66.8$\pm$0.1 66.1$\pm$0.3 68.4$\pm$0.3 70.0$\pm$0.1 69.8$\pm$0.2 70.1$\pm$0.3
Inductive Flickr SGC 44 0.1% 27.6$\pm$0.1 34.5$\pm$0.1 43.7$\pm$0.5 43.7$\pm$0.2 41.8$\pm$0.4 42.2$\pm$0.1 43.6$\pm$0.3 40.3$\pm$0.0 37.5$\pm$0.2 45.3$\pm$0.2 45.3$\pm$0.7 44.2±0.0
223 0.5% 33.5$\pm$0.0 36.1$\pm$0.0 42.2$\pm$0.2 43.8$\pm$0.6 42.6$\pm$0.2 43.9$\pm$0.1 41.6$\pm$1.6 40.8$\pm$0.1 36.2$\pm$0.2 43.6$\pm$0.1 45.7$\pm$0.4
446 1% 34.4$\pm$0.0 36.5$\pm$0.0 41.1$\pm$0.8 43.7$\pm$0.3 42.1$\pm$0.4 43.3$\pm$0.1 43.2$\pm$0.4 42.7$\pm$0.4 34.3$\pm$0.3 42.6$\pm$0.1 46.1$\pm$0.5
GCN 44 0.1% 39.2$\pm$0.0 40.7$\pm$0.0 43.3$\pm$0.3 45.4$\pm$0.3 43.3$\pm$1.1 44.9$\pm$0.9 42.4$\pm$0.2 44.5$\pm$0.4 44.8$\pm$0.5 45.2$\pm$0.3 46.9$\pm$0.3 46.8±0.2
223 0.5% 42.4$\pm$0.0 41.4$\pm$0.0 44.6$\pm$0.4 45.8$\pm$0.3 43.5$\pm$1.0 45.0$\pm$0.2 44.9$\pm$0.3 45.0$\pm$0.2 46.4$\pm$0.2 45.6$\pm$0.7 47.0$\pm$0.1
446 1% 43.4$\pm$0.0 41.4$\pm$0.0 44.4$\pm$0.1 45.7$\pm$0.2 42.8$\pm$0.2 44.8$\pm$0.4 45.2$\pm$0.2 44.6$\pm$0.3 46.7$\pm$0.2 44.9$\pm$0.1 47.2$\pm$0.1
Reddit SGC 153 0.05% 55.8$\pm$0.2 54.0$\pm$0.1 89.7$\pm$0.6 91.1$\pm$0.1 90.8$\pm$0.1 91.0$\pm$0.0 90.5$\pm$0.3 90.3$\pm$0.8 92.0$\pm$0.0 92.1$\pm$0.0 90.9$\pm$0.2 93.2±0.0
769 0.1% 74.1$\pm$0.1 78.6$\pm$0.0 91.8$\pm$0.2 90.4$\pm$0.1 91.5$\pm$0.0 91.8$\pm$0.0 91.9$\pm$0.0 88.1$\pm$2.8 90.9$\pm$0.1 91.0$\pm$0.0 92.6$\pm$0.2
1539 0.2% 83.3$\pm$0.0 83.8$\pm$0.0 92.1$\pm$0.3 90.4$\pm$0.1 92.0$\pm$0.2 92.1$\pm$0.0 86.3$\pm$5.6 91.7$\pm$0.2 90.0$\pm$0.0 89.6$\pm$0.1 92.6$\pm$0.3
GCN 153 0.05% 60.4$\pm$0.0 58.6$\pm$0.1 56.8$\pm$2.1 85.8$\pm$0.1 75.8$\pm$1.9 88.6$\pm$0.3 72.9$\pm$4.9 88.9$\pm$1.2 91.2$\pm$0.1 91.3$\pm$0.1 89.2$\pm$0.5 94.2±0.0
769 0.1% 81.7$\pm$0.0 81.7$\pm$0.0 87.4$\pm$0.4 90.5$\pm$0.2 87.9$\pm$1.1 91.2$\pm$0.2 89.6$\pm$2.5 91.8$\pm$0.3 92.4$\pm$0.0 92.6$\pm$0.2 90.9$\pm$0.3
1539 0.2% 87.1$\pm$0.0 86.9$\pm$0.0 91.4$\pm$0.4 89.5$\pm$0.3 88.2$\pm$2.1 91.9$\pm$0.1 91.2$\pm$0.3 92.2$\pm$0.1 92.7$\pm$0.1 92.5$\pm$0.3 92.4$\pm$0.1

### v3

@@ 2 Preliminary
Eigenbasis matching [18] methods align condensed graphs and original graphs in the spectral domain via eigenbasis:
$\mathcal{L}_{e}=\sum_{k=1}^{K}\left\|\boldsymbol{X}^{\top}\boldsymbol{u}_{k}\boldsymbol{u}_{k}^{\top}\boldsymbol{X}-\boldsymbol{X}^{\prime\top}\boldsymbol{u}_{k}^{\prime}\boldsymbol{u}_{k}^{\prime\top}\boldsymbol{X}^{\prime}\right\|_{F}^{2},$ (6)
where $\boldsymbol{u}_{k}\boldsymbol{u}_{k}^{\top}$ and $\boldsymbol{u}_{k}^{\prime}\boldsymbol{u}_{k}^{\prime\top}$ are the subspaces induced by the $k$-th eigenvector in the real and synthetic graphs. A discrimination constraint to preserve the category-level information:
$\mathcal{L}_{d}=\sum_{i=1}^{C}\left(1-\frac{\boldsymbol{H}_{i}^{\top}\cdot\boldsymbol{H}^{\prime}_{i}}{\|\boldsymbol{H}_{i}\|\|\boldsymbol{H}^{\prime}_{i}\|}\right)+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{C}\frac{\boldsymbol{H}_{i}^{\top}\cdot\boldsymbol{H}^{\prime}_{j}}{\|\boldsymbol{H}_{i}\|\|\boldsymbol{H}^{\prime}_{j}\|},$ (7)
where $\boldsymbol{H}=\boldsymbol{Y}^{\top}\boldsymbol{A}\boldsymbol{X}$ and $\boldsymbol{H}^{\prime}=\boldsymbol{Y}^{\prime\top}\sum_{k=1}^{K}(1-\lambda_{k})\boldsymbol{u}_{k}^{\prime}\boldsymbol{u}_{k}^{\prime\top}$. An additional regularisation is used to constrain the representation space:
$\mathcal{L}_{o}=\left\|\boldsymbol{U}_{K}^{\prime\top}\boldsymbol{U}_{K}^{\prime}-\boldsymbol{I}_{K}\right\|_{F}^{2}.$ (8)
The overall matching loss function of GDEM is formulated as the weighted sum of three regularisation terms:
$\mathcal{L}_{\text{match}}=\alpha\mathcal{L}_{e}+\beta\mathcal{L}_{d}+\gamma\mathcal{L}_{o}.$ (9)
@@ 4.1 Baselines
Trajectory matching: SFGC [36] firstly generates training trajectories of training a GNN backbone using the original graph. For the condensation phase, a GNN with the same architecture is trained with the condensed graph, and the model parameters are matched to the generated model parameters. GEOM [32] applies graph curriculum learning in the trajectory generation phase and uses expanding window matching and knowledge distillation to improve the quality of the condensed graph in the condensation phase.
Eigenbasis matching: GDEM [18] first performs eigendecomposition on the original dataset, then matches the synthetic bases with the real ones to reflect the structure and feature relations.
@@ 4.2 Implementations
Table 2: Overall node classification performance. For each dataset with different condensed graph sizes and backbone GNN models, the red and bold results indicate the best performance, while the blue and italic results indicate the runner-up. OOT stands for ‘Out of Time’ for 48 hours (e.g., GDEM runs out of time during the eigendecomposition phase).
@@
Dataset GNN Size Ratio Traditional Gradient Distribution Trajectory Eigenbasis Whole
@@
Random k-Center GCond GCondX DosCond DosCondX SGDD GCDM GCDMX DM SFGC GEOM GDEM
Transductive CiteSeer SGC 30 0.9% 53.8$\pm$0.1 52.7$\pm$0.0 71.9$\pm$0.6 65.3$\pm$0.1 71.1$\pm$0.9 64.3$\pm$0.4 71.1$\pm$0.1 66.0$\pm$2.2 70.1$\pm$0.8 64.2$\pm$8.5 65.2$\pm$0.3 60.1$\pm$0.2 70.8$\pm$0.5 70.3±1.0
60 1.8% 61.9$\pm$0.0 66.8$\pm$0.0 71.0$\pm$0.6 70.9$\pm$0.3 72.0$\pm$1.1 70.6$\pm$0.1 69.9$\pm$0.1 66.7$\pm$0.0 65.8$\pm$1.3 65.7$\pm$2.5 67.0$\pm$0.8 65.2$\pm$0.2 72.3$\pm$0.6
120 3.6% 68.1$\pm$0.0 68.1$\pm$0.0 72.5$\pm$1.2 71.5$\pm$0.0 70.6$\pm$0.2 71.8$\pm$0.3 70.8$\pm$0.8 69.1$\pm$1.2 68.1$\pm$0.2 69.7$\pm$0.1 68.8$\pm$0.2 67.7$\pm$0.3 72.8$\pm$0.4
GCN 30 0.9% 58.9$\pm$0.0 65.0$\pm$0.0 46.3$\pm$7.0 66.7$\pm$4.2 69.2$\pm$0.5 67.7$\pm$1.2 70.6$\pm$1.5 71.2$\pm$0.8 72.8$\pm$0.3 72.6$\pm$0.6 69.7$\pm$0.3 69.6$\pm$0.6 71.7$\pm$0.3 71.4±0.5
60 1.8% 62.6$\pm$0.0 67.8$\pm$0.0 54.2$\pm$3.9 69.6$\pm$0.4 70.4$\pm$1.7 70.1$\pm$0.2 71.5$\pm$0.7 71.9$\pm$0.7 71.7$\pm$0.2 72.2$\pm$0.7 69.4$\pm$0.0 67.5$\pm$0.9 72.7$\pm$0.6
120 3.6% 69.6$\pm$0.0 69.4$\pm$0.0 70.7$\pm$0.7 70.9$\pm$0.2 47.3$\pm$7.3 70.4$\pm$0.2 71.0$\pm$0.7 72.3$\pm$1.3 72.5$\pm$0.5 72.4$\pm$0.1 69.8$\pm$0.5 72.1$\pm$1.0 73.4$\pm$0.4
Cora SGC 35 1.3% 59.3$\pm$0.0 63.8$\pm$0.0 80.6$\pm$0.1 80.6$\pm$0.2 80.6$\pm$0.1 80.8$\pm$0.1 62.4$\pm$5.5 77.0$\pm$0.4 79.9$\pm$0.1 77.0$\pm$0.5 73.8$\pm$1.5 69.2$\pm$1.2 71.0$\pm$0.6 79.2±0.6
70 2.6% 70.7$\pm$0.0 70.3$\pm$0.0 81.0$\pm$0.2 79.0$\pm$0.3 80.3$\pm$0.5 79.8$\pm$0.2 80.8$\pm$0.4 78.9$\pm$1.0 79.7$\pm$0.3 78.0$\pm$1.6 77.5$\pm$0.1 69.6$\pm$1.5 75.4$\pm$0.5
140 5.2% 77.0$\pm$0.0 77.1$\pm$0.0 80.9$\pm$0.4 81.2$\pm$0.3 80.8$\pm$0.4 80.8$\pm$0.5 81.4$\pm$0.4 77.9$\pm$0.7 80.1$\pm$0.2 79.6$\pm$0.6 79.2$\pm$0.1 77.3$\pm$0.1 79.5$\pm$0.8
GCN 35 1.3% 63.9$\pm$0.1 66.5$\pm$0.0 80.5$\pm$0.4 79.8$\pm$1.4 72.0$\pm$8.7 80.1$\pm$0.9 80.5$\pm$0.4 78.9$\pm$0.8 79.1$\pm$0.9 79.2$\pm$0.2 79.6$\pm$0.2 80.3$\pm$1.1 68.0$\pm$0.1 81.7±0.9
70 2.6% 73.0$\pm$0.0 71.6$\pm$0.0 78.1$\pm$3.6 80.6$\pm$0.9 79.6$\pm$0.7 81.1$\pm$0.3 81.2$\pm$0.6 79.4$\pm$0.6 80.5$\pm$0.3 79.6$\pm$0.3 79.5$\pm$0.1 81.5$\pm$0.8 72.8$\pm$0.8
140 5.2% 77.1$\pm$0.0 76.6$\pm$0.0 80.2$\pm$1.7 81.5$\pm$0.1 80.5$\pm$0.7 80.6$\pm$0.2 79.9$\pm$1.6 79.9$\pm$0.2 80.2$\pm$0.5 79.9$\pm$0.3 80.1$\pm$0.6 82.2$\pm$0.4 77.4$\pm$0.6
PubMed SGC 15 0.08% 67.8$\pm$0.2 70.5$\pm$0.1 75.9$\pm$0.7 77.3$\pm$0.2 74.2$\pm$1.1 75.7$\pm$0.5 76.4$\pm$0.9 73.3$\pm$1.2 74.0$\pm$0.3 72.1$\pm$0.9 73.9$\pm$0.5 73.8$\pm$0.3 73.8$\pm$0.6 76.9±0.1
30 0.15% 72.5$\pm$0.2 75.8$\pm$0.0 75.2$\pm$0.0 77.1$\pm$0.2 77.2$\pm$0.1 77.0$\pm$0.1 78.0$\pm$0.3 74.7$\pm$0.6 75.2$\pm$0.7 75.1$\pm$0.6 75.8$\pm$0.2 77.4$\pm$0.4 78.7$\pm$0.4
60 0.3% 75.6$\pm$0.0 75.7$\pm$0.0 75.7$\pm$0.0 76.8$\pm$0.1 75.5$\pm$0.2 75.5$\pm$0.0 76.1$\pm$0.1 76.5$\pm$1.1 76.3$\pm$0.2 75.8$\pm$0.0 75.8$\pm$0.0 75.8$\pm$0.4 79.1$\pm$0.1
GCN 15 0.08% 69.8$\pm$0.1 72.1$\pm$0.1 67.6$\pm$10.4 77.8$\pm$0.3 75.7$\pm$0.7 75.7$\pm$0.1 76.7$\pm$1.1 75.9$\pm$0.6 77.1$\pm$0.3 76.0$\pm$0.7 78.4$\pm$0.1 80.1$\pm$0.3 73.3$\pm$0.6 79.3±0.3
30 0.15% 73.7$\pm$0.1 76.4$\pm$0.0 74.6$\pm$0.8 78.0$\pm$0.5 76.8$\pm$0.2 78.6$\pm$0.2 78.5$\pm$0.4 77.4$\pm$0.4 76.8$\pm$0.6 77.5$\pm$0.1 78.1$\pm$0.4 79.7$\pm$0.3 78.3$\pm$0.8
60 0.3% 78.0$\pm$0.0 78.2$\pm$0.0 77.2$\pm$0.7 78.0$\pm$0.1 77.3$\pm$1.2 78.0$\pm$0.1 78.0$\pm$1.1 77.6$\pm$0.4 78.1$\pm$0.3 78.0$\pm$0.2 78.5$\pm$0.5 79.5$\pm$0.4 78.8$\pm$0.3
Arxiv SGC 90 0.05% 45.6$\pm$0.2 51.8$\pm$0.2 65.5$\pm$0.0 66.0$\pm$0.2 62.7$\pm$0.6 61.6$\pm$0.3 64.5$\pm$0.9 60.8$\pm$0.1 59.8$\pm$0.3 61.0$\pm$0.2 66.1$\pm$0.2 62.0$\pm$0.5 OOT 68.8±0.0
454 0.25% 55.2$\pm$0.0 58.2$\pm$0.0 66.5$\pm$0.5 66.4$\pm$0.1 63.7$\pm$0.2 63.8$\pm$0.1 66.4$\pm$0.3 62.7$\pm$0.9 61.8$\pm$0.5 62.9$\pm$0.2 66.7$\pm$0.3 62.8$\pm$0.7 OOT
909 0.5% 58.3$\pm$0.0 60.3$\pm$0.0 67.2$\pm$0.1 67.4$\pm$0.3 63.9$\pm$0.1 64.3$\pm$0.4 66.9$\pm$0.3 62.4$\pm$0.2 62.6$\pm$0.2 62.5$\pm$0.0 66.4$\pm$0.3 63.6$\pm$0.3 OOT
GCN 90 0.05% 47.1$\pm$0.0 54.5$\pm$0.0 53.7$\pm$1.6 62.7$\pm$0.4 55.6$\pm$0.4 61.0$\pm$0.5 55.9$\pm$5.8 63.3$\pm$0.3 63.8$\pm$0.3 64.4$\pm$0.5 66.1$\pm$0.4 65.5$\pm$1.0 OOT 71.1±0.0
454 0.25% 56.8$\pm$0.0 60.3$\pm$0.0 64.2$\pm$0.2 65.4$\pm$0.4 61.6$\pm$0.5 64.7$\pm$0.2 63.2$\pm$0.3 66.4$\pm$0.1 66.7$\pm$0.4 67.5$\pm$0.3 67.2$\pm$0.4 65.8$\pm$0.4 OOT
909 0.5% 60.3$\pm$0.0 62.1$\pm$0.0 65.1$\pm$0.4 66.6$\pm$0.2 63.4$\pm$0.4 65.8$\pm$0.1 66.8$\pm$0.3 67.6$\pm$0.0 67.6$\pm$0.3 68.0$\pm$0.3 67.8$\pm$0.2 66.2$\pm$0.5 OOT
Products SGC 612 0.025% 51.6$\pm$1.3 48.6$\pm$0.6 64.0$\pm$0.2 64.6$\pm$0.1 62.1$\pm$0.1 63.6$\pm$0.3 64.9$\pm$0.1 57.7$\pm$0.2 58.9$\pm$0.1 58.5$\pm$0.6 62.2$\pm$0.1 61.1$\pm$0.4 OOT 64.7±0.1
1225 0.05% 55.2$\pm$0.8 52.2$\pm$0.7 64.0$\pm$0.1 62.4$\pm$0.1 61.0$\pm$0.3 62.7$\pm$0.4 62.3$\pm$0.2 58.2$\pm$0.3 61.0$\pm$0.1 60.9$\pm$0.1 62.2$\pm$0.2 62.4$\pm$0.2 OOT
2449 0.1% 58.0$\pm$0.8 55.4$\pm$0.4 64.4$\pm$0.4 62.8$\pm$0.1 61.4$\pm$0.2 62.3$\pm$0.2 64.3$\pm$0.3 60.8$\pm$0.2 61.3$\pm$0.1 61.3$\pm$0.2 61.9$\pm$0.2 63.1$\pm$0.2 OOT
GCN 612 0.025% 57.8$\pm$0.5 55.4$\pm$0.8 63.7$\pm$0.3 65.5$\pm$0.2 61.2$\pm$0.3 62.2$\pm$0.6 64.0$\pm$0.4 66.5$\pm$0.1 68.0$\pm$0.3 66.2$\pm$0.1 67.1$\pm$0.2 68.5$\pm$0.3 OOT 73.1±0.1
1225 0.05% 61.6$\pm$0.5 57.6$\pm$0.7 67.0$\pm$0.2 66.6$\pm$0.3 62.6$\pm$0.4 64.2$\pm$0.2 65.9$\pm$0.2 68.4$\pm$0.4 68.8$\pm$0.1 68.5$\pm$0.2 67.9$\pm$0.3 69.8$\pm$0.3 OOT
2449 0.1% 65.3$\pm$0.5 59.1$\pm$0.5 68.0$\pm$0.2 68.3$\pm$0.2 65.8$\pm$0.2 66.8$\pm$0.1 66.1$\pm$0.3 68.4$\pm$0.3 70.0$\pm$0.1 69.8$\pm$0.2 70.1$\pm$0.3 71.1$\pm$0.3 OOT
Inductive Flickr SGC 44 0.1% 27.6$\pm$0.1 34.5$\pm$0.1 43.7$\pm$0.5 43.7$\pm$0.2 41.8$\pm$0.4 42.2$\pm$0.1 43.6$\pm$0.3 40.3$\pm$0.0 37.5$\pm$0.2 45.3$\pm$0.2 45.3$\pm$0.7 33.6$\pm$0.4 OOT 44.2±0.0
223 0.5% 33.5$\pm$0.0 36.1$\pm$0.0 42.2$\pm$0.2 43.8$\pm$0.6 42.6$\pm$0.2 43.9$\pm$0.1 41.6$\pm$1.6 40.8$\pm$0.1 36.2$\pm$0.2 43.6$\pm$0.1 45.7$\pm$0.4 37.4$\pm$0.2 OOT
446 1% 34.4$\pm$0.0 36.5$\pm$0.0 41.1$\pm$0.8 43.7$\pm$0.3 42.1$\pm$0.4 43.3$\pm$0.1 43.2$\pm$0.4 42.7$\pm$0.4 34.3$\pm$0.3 42.6$\pm$0.1 46.1$\pm$0.5 38.1$\pm$0.2 OOT
GCN 44 0.1% 39.2$\pm$0.0 40.7$\pm$0.0 43.3$\pm$0.3 45.4$\pm$0.3 43.3$\pm$1.1 44.9$\pm$0.9 42.4$\pm$0.2 44.5$\pm$0.4 44.8$\pm$0.5 45.2$\pm$0.3 46.9$\pm$0.3 44.6$\pm$0.5 OOT 46.8±0.2
223 0.5% 42.4$\pm$0.0 41.4$\pm$0.0 44.6$\pm$0.4 45.8$\pm$0.3 43.5$\pm$1.0 45.0$\pm$0.2 44.9$\pm$0.3 45.0$\pm$0.2 46.4$\pm$0.2 45.6$\pm$0.7 47.0$\pm$0.1 45.2$\pm$0.9 OOT
446 1% 43.4$\pm$0.0 41.4$\pm$0.0 44.4$\pm$0.1 45.7$\pm$0.2 42.8$\pm$0.2 44.8$\pm$0.4 45.2$\pm$0.2 44.6$\pm$0.3 46.7$\pm$0.2 44.9$\pm$0.1 47.2$\pm$0.1 45.5$\pm$0.1 OOT
Reddit SGC 153 0.05% 55.8$\pm$0.2 54.0$\pm$0.1 89.7$\pm$0.6 91.1$\pm$0.1 90.8$\pm$0.1 91.0$\pm$0.0 90.5$\pm$0.3 90.3$\pm$0.8 92.0$\pm$0.0 92.1$\pm$0.0 90.9$\pm$0.2 59.4$\pm$1.5 OOT 93.2±0.0
769 0.1% 74.1$\pm$0.1 78.6$\pm$0.0 91.8$\pm$0.2 90.4$\pm$0.1 91.5$\pm$0.0 91.8$\pm$0.0 91.9$\pm$0.0 88.1$\pm$2.8 90.9$\pm$0.1 91.0$\pm$0.0 92.6$\pm$0.2 81.7$\pm$0.7 OOT
1539 0.2% 83.3$\pm$0.0 83.8$\pm$0.0 92.1$\pm$0.3 90.4$\pm$0.1 92.0$\pm$0.2 92.1$\pm$0.0 86.3$\pm$5.6 91.7$\pm$0.2 90.0$\pm$0.0 89.6$\pm$0.1 92.6$\pm$0.3 86.7$\pm$0.1 OOT
GCN 153 0.05% 60.4$\pm$0.0 58.6$\pm$0.1 56.8$\pm$2.1 85.8$\pm$0.1 75.8$\pm$1.9 88.6$\pm$0.3 72.9$\pm$4.9 88.9$\pm$1.2 91.2$\pm$0.1 91.3$\pm$0.1 89.2$\pm$0.5 90.0$\pm$0.5 OOT 94.2±0.0
769 0.1% 81.7$\pm$0.0 81.7$\pm$0.0 87.4$\pm$0.4 90.5$\pm$0.2 87.9$\pm$1.1 91.2$\pm$0.2 89.6$\pm$2.5 91.8$\pm$0.3 92.4$\pm$0.0 92.6$\pm$0.2 90.9$\pm$0.3 89.4$\pm$0.5 OOT
1539 0.2% 87.1$\pm$0.0 86.9$\pm$0.0 91.4$\pm$0.4 89.5$\pm$0.3 88.2$\pm$2.1 91.9$\pm$0.1 91.2$\pm$0.3 92.2$\pm$0.1 92.7$\pm$0.1 92.5$\pm$0.3 92.4$\pm$0.1 91.2$\pm$0.1 OOT